use newton raphson method :


sin ( x + π/2) - ln |x| =0


1
Expert's answer
2022-05-12T09:29:02-0400

We remind the Newton-Raphson method. Suppose that "f(x)" is a differentiable function on the interval "(t_0,t_1)", "t_0<t_1", "r\in(t_0,t_1)" is a root of equation "f(x)=0" and "x_0=r+h\in(t_0,t_1)" is a point near the root "r". Define: "x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}", "n\in{\mathbb{N}}". Then, the sequence "x_n" converges to the root "r". The error after "n+1" iterations is proportional to the square of error after "n" iterations. The method works in case there are no other roots in the neighborhood of "r".

Consider the function "f(x)=\sin(x+\frac{\pi}2)-\ln(|x|)=\cos(x)-\ln(|x|)". For "|x|>e" there are no roots, since "ln(|x|)>1" and "|\sin(x+\frac{\pi}{2})|\leq1". Analysis of the graph of the function "f(x)" with the help of mathematical packages shows one root at the interval "(-e,0)" and one root at the interval "(0,e)". Consider the function for "x>0". We get: "f(x)=\cos(x)-\ln(x)". "f'(x)=-\sin(x)-\frac1x". "f'<0" for "x\in(0,1)". Set "x_0=1" and perform the method:

"x_1=x_0-\frac{f(x_0)}{f'(x_0)}\approx1.29341", "x_2=x_1-\frac{f(x_1)}{f'(x_1)}\approx1.30296", "x_3=x_2-\frac{f(x_2)}{f'(x_2)}\approx1.30296".

The method can be stopped at "x_4". The root is: "r_1\approx1.30296." All values are rounded up to "5" decimal places. Since the function is even, another root is: "r_2\approx-1.30296."

Answer: the roots are "r_1\approx1.30296" and "r_2\approx-1.30296." The values are rounded to "5" decimal places.